# MovieDreamer: Hierarchical Generation for Coherent Long Visual Sequences

### v2

@@ Abstract
Recent advancements in video generation have primarily leveraged diffusion models for short-duration content. However, these approaches often fall short in modeling complex narratives and maintaining character consistency over extended periods, which is essential for long-form video production like movies. We propose MovieDreamer, a novel hierarchical framework that integrates the strengths of autoregressive models with diffusion-based rendering to pioneer long-duration video generation with intricate plot progressions and high visual fidelity. Our approach utilizes autoregressive models for global narrative coherence, predicting sequences of visual tokens that are subsequently transformed into high-quality video frames through diffusion rendering. This method is akin to traditional movie production processes, where complex stories are factorized down into manageable scene capturing. Further, we employ a multimodal script that enriches scene descriptions with detailed character information and visual style, enhancing continuity and character identity across scenes. We present extensive experiments across various movie genres, demonstrating that our approach not only achieves superior visual and narrative quality but also effectively extends the duration of generated content significantly beyond current capabilities. Homepage: https://aim-uofa.github.io/MovieDreamer/.
@@ 1 Introduction
Driven by the advent of generative modeling techniques (Rombach et al., 2022; Podell et al., 2023), there has been significant progress in video generation in the research community (Guo et al., 2024; Ho et al., 2022b; Blattmann et al., 2023). Tremendous research efforts have focused on adapting a pre-trained text-to-image diffusion model (Guo et al., 2024; Ho et al., 2022b) to a video generation model. Recently revolutionary leap has been made with the Sora (Brooks et al., 2024) model which substantially scales a spatial-temporal transformer model. This model demonstrates remarkable video generation quality, dramatically extending the duration of generated videos from seconds to an unprecedented full minute. This technology profoundly reshapes the boundaries of generative AI, fostering anticipation towards hours-long movie production.
While predominant video generation methods adopt diffusion models, this paradigm is more suited for visual rendering and is less adept at modeling complex abstract logic and reasoning compared to autoregressive models, as evidenced in natural language processing. Moreover, diffusion models lack the flexibility to support arbitrary length. On the other hand, autoregressive models (Wang et al., 2023a; Jiang et al., 2022; Hu et al., 2023; Pan et al., 2023) have shown superior ability in handling complex reasoning and are better at predicting what might happen in the next—a key element of a world model. Moreover, autoregressive models offer greater flexibility in handling varied lengths of predictions and can also benefit from mature training and inference infrastructure. As such, there are attempts that use the autoregressive model for video generation (Kondratyuk et al., 2023; Hu et al., 2023). However, autoregressive modeling is not as compute efficient as diffusion model for visual rendering and demands substantially more computing even for image rendering.
@@
Recent progress (Henschel et al., 2024; Chen et al., 2023; Qiu et al., 2023; Brooks et al., 2024) suggests that video generation quality in short duration (from seconds to a minute) can be consistently improved by scaling the compute due to the scaling law. In this work, we sidestep the short video generation problem and instead seek a generative framework for long video generation featuring complex narrative structures and intricate plot progressions, which is hardly achieved by merely scaling the current approaches. Our aim is to push the boundaries of video generation beyond the limitations of short-duration content, enabling the creation of videos with rich, engaging storylines.
@@
To this end, we represent the whole movie using sparse keyframes and employ a diffusion autoencoder to tokenize each keyframe into compact visual tokens. We train an autoregressive model to predict the sequence of visual tokens. This model is initialized from a pretrained large language mode, a 7B LLaMA model (Touvron et al., 2023), to leverage its rich world knowledge for better generalizability. As such, the model can be regarded as a multimodal model, which, rather than outputting the text, predicts the visual tokens with movie script conditioning.
@@ 2 Related Work
Video generation has advanced mainly through diffusion and autoregression methods. The popularization of Stable Diffusion (Rombach et al., 2022) has sparked extensive studies, improving video quality with various techniques (Guo et al., 2024; Ho et al., 2022b; a; He et al., 2023). For long videos, strategies like (Yin et al., 2023; Qiu et al., 2023; Wang et al., 2023b; Chen et al., 2023; Henschel et al., 2024) can only generate simple results. Despite their high computational load and challenges in managing complex sequences, diffusion models are widely used. Autoregression, on the other hand, is less common but excels in specific areas like autonomous driving by employing unified token spaces (Hu et al., 2023), although it struggles with continuity in complex scenes. Our approach combines the strengths of both methods in a hierarchical and scalable manner, enhancing control and flexibility irrespective of video length.
@@
Recent advancements in NLP lead to the emergence of visual language models (VLMs) through enabling large language model (Touvron et al., 2023) to comprehend visual modalities (Liu et al., 2023b; Bai et al., 2023a; b; Chen et al., 2024). In addition to language, VLMs have demonstrated their power in generating images (Pan et al., 2023; Sun et al., 2023c; b). Moreover, ongoing research (Zhang et al., 2023; Lin et al., 2023; Li et al., 2023a; Zhao et al., 2022) has showcased the robust capabilities of Multimodal Language Models (MLLMs) in comprehending video content. Some works (Cai et al., 2024; Cheng et al., 2024; Ren et al., 2024b; Rasheed et al., 2024; Lai et al., 2024) enable VLMs to assist in spatial understanding and other visual tasks. VLMs (Jiang et al., 2022; Kim et al., 2024; Etukuru et al., 2024; Brohan et al., 2023; Driess et al., 2023) also demonstrate the powerful capabilities in robotics. These advancements have solidified the effectiveness of VLMs in addressing complex problems. Building on this foundation, our approach leverages VLMs to facilitate the generation of extremely long video content.
@@
The crux of visual story generation lies in ensuring consistency across generated images. Similar to video generation, diffusion and autoregression are also two primary manners. Existing diffusion-based methods predominantly employ conditional generation to maintain the coherence  (Su et al., 2023; Wang et al., 2023c; Tewel et al., 2024; Hertz et al., 2023). Some (Li et al., 2023b; Wu et al., 2024) have shown promising results in retaining high-quality facial features and identity. Other approaches are based on the autoregressive ideas  (Feng et al., 2023; Liu et al., 2023a; Rahman et al., 2023; Pan et al., 2022). However, these approaches generally yield mediocre results and are challenging to scale up. Story and video generation share intrinsic links, with overlapping solutions. Our method uniquely integrates these, offering a unified approach.
@@ 3.2 Keyframe Tokenization via Diffusion Autoencoder
To create a concise yet faithful representation of images, we employ a diffusion autoencoder. Our encoder, $\mathcal{E}$, contains a pretrained CLIP vision model (Radford et al., 2021) and a transformer-based token compressor, which encodes an image $\mathbf{x}$ into a reduced number of compressed tokens denoted by $\mathbf{e}$. The decoder, $\mathcal{D}$, is finetuned based on the pretrained SDXL and yields an 896$\times$512 reconstructed image by leveraging the compressed tokens in the cross-attention module. We train the compressor and the decoder while leaving the CLIP vision model frozen. The loss for training this diffusion autoencoder is as follows:
@@
where $\mathbf{x}_{0}$ is the input image and $\mathbf{z}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\bm{\epsilon}$ is its corresponding noisy latent at timestep $t$. $\bm{\epsilon}\in\mathcal{N}(0,I)$ is the Gaussian noise. $\alpha_{t}$ and $\sigma_{t}$ define the noise schedule. In our experiment, we empirically find that merely two tokens can sufficiently characterize major semantics of keyframes, corroborating the finding of previous work (Yang et al., 2022).
@@ 3.3 ID-preserving Diffusion Rendering
In our approach, while the primary diffusion decoder $\mathcal{D}$ adeptly reconstructs target images, it occasionally falls short in capturing fine-grained details, notably in facial features, due to the attenuation of details in compressed tokens. To address this, we enhance the cross-attention module within $\mathcal{D}$. This enhancement involves the integration of both descriptive text embedding, $\mathbf{d}$, and face embedding, $\mathbf{f}$, derived from the multimodal script. Specifically, the face embedding and description embedding is projected to the same dimension as the compressed token with two additional MLPs, which are then concatenated together and serve as the input to the keys and values of cross-attention modules in $\mathcal{D}$.
To further advocate the model’s ability to focus on critical details, we introduce a random masking strategy that obscures a subset of the input tokens. This technique encourages the decoder to more effectively utilize the available facial and textual cues to reconstruct the images with higher fidelity, particularly in preserving identity-specific characteristics. This ID-preserving rendering also compensates for the loss of identity during the autoregressive modeling, leading to substantially improved identity perception quality as illustrated in Figure 3.
@@ 3.4 Autoregressive Keyframe Token Generation
We initiate our approach by leveraging a pretrained large language model, specifically LLaMA2-7B (Touvron et al., 2023), to construct our autoregressive model, $\mathcal{G}$. Unlike traditional LLMs, $\mathcal{G}$ is designed to predict compressed vision tokens from the multimodal script $\mathbf{M}$ of the corresponding keyframe and historical information $\mathbf{H}$, which is formulated as $\mathbf{e}_{t}=\mathcal{G}(\mathbf{H}_{<t},\mathbf{M}_{t})$.
Traditional LLMs often employ cross-entropy loss for training, which is suitable for discrete outputs. However, our model deals with continuous real-valued image tokens, which makes cross-entropy inapplicable. Inspired by  Tschannen et al. (2024), we adopt a $k$-mixture Gaussian Mixture Model (GMM) to effectively model the distribution of these real-valued tokens. This involves parameterizing the GMM with $2kd$ means, $2kd$ variances, and $k$ mixing coefficients. These parameters are predicted by attaching an additional trainable linear layer at the end of the autoregressive model, enabling the construction of the GMM and the sampling of continuous tokens from the GMM. The model is trained by minimizing the negative log-likelihood:
@@
We develop a well-structured multi-modal script format to serve as input for autoregressive models, as depicted in Figure 16. Our script integrates multiple dimensions: characters, scene elements, and plots. Accurately representing character appearance using text alone is challenging; therefore, we combine textual descriptions with facial embeddings (Zheng et al., 2022) to provide a more detailed representation of each character. To facilitate the processing by the autoregressive model, we structure the script format to distinctly delineate these elements.
@@
This method significantly extends the usable contextual length during training by condensing an entire sentence into a single token. We utilize LongCLIP (Zhang et al., 2024) as our text encoder for descriptions, supporting inputs up to 248 tokens, which enhances our capacity to handle detailed narrative content. Consequently, the multi-modal script at timestep $t$ and preceding historical information are represented as:
@@
To promote personalized movie content generation, we propose a few-shot learning method that utilizes in-context learning. During training, we select 10 random frames from an episode, encode them into visual tokens, and prepend these tokens stochastically to the episode’s visual tokens. This strategy not only promotes in-context learning, allowing the model to tailor content based on the reference frames, but also acts as a data augmentation technique, effectively mitigating overfitting.
@@ 3.5 Keyframe-based Video Generation
After obtaining the keyframes, we can generate video clips of the movie based on these keyframes. A straightforward approach is to utilize an existing image-to-video model, such as Stable Video Diffusion (SVD) (Blattmann et al., 2023), to generate these clips. Specifically, SVD transforms the input image into latent features for conditioning and introduces interaction with the CLIP features of the input image via cross-attention. Although SVD is capable of generating high-quality short videos, e.g., 25 frames, it struggles to generate longer movie clips.
@@ 4.1 Implementation Details
We use pretrained models to reduce the training computational overhead, which is very common in the era of large models. We use the U-Net of pretrained SDXL (Podell et al., 2023) as our diffusion decoder. The encoder consists of a pretrained CLIP image encoder (ViT-L) (Radford et al., 2021) and a transformer-based compressor. The autoregressive model is initialized with a pretrained LLaMA-7B (Touvron et al., 2023) checkpoint. All MLPs used to align different modal inputs follow the architecture in LLaVA (Liu et al., 2023b). Please refer to Appendix A for more details about the network architecture and training.
@@ 4.2 Experimental Setup
The majority of our training data is sourced from Condensed Movies (Bain et al., 2020), with additional data collected using the same methodology as this dataset, resulting in 5M keyframes with corresponding script annotations. To systematically evaluate the effectiveness of our method, we construct a test dataset consisting of 100 long movies that are NOT included in the training set, with 1M keyframes after pre-processing. Each of these videos is processed through our proposed data pre-processing pipeline (see Appendix C), which accurately extracts and annotates keyframes.
@@
We evaluate the effectiveness of various methods using the following metrics. The CLIP score (Radford et al., 2021) assesses semantic alignment between the plot and generated outputs. Quality is evaluated using the Aesthetic Score (AS) (Schuhmann et al., 2022), Fréchet Image Distance (FID) (Heusel et al., 2018), and Inception Score (IS) (Salimans et al., 2016). For assessing consistency, we propose new metrics for short-term (ST) and long-term (LT) consistency. ST is measured by calculating the CLIP cosine similarity between adjacent frames containing the target character. LT consistency is evaluated by checking the presence of the target character in selected frames across the test set. Finally, we calculate the ratio of frames containing the target character to the total number of selected frames as a measure of the accuracy in maintaining long-term character consistency. Please refer to the Appendix B for more details about the evaluation metrics.

### v3

@@ Abstract
Recent advancements in video generation have primarily leveraged diffusion models for short-duration content. However, these approaches often fall short in modeling complex narratives and maintaining character consistency over extended periods, which is essential for long-form video production like movies. We propose MovieDreamer, a novel hierarchical framework that integrates the strengths of autoregressive models with diffusion-based rendering to pioneer long-duration video generation with intricate plot progressions and high visual fidelity. Our approach utilizes autoregressive models for global narrative coherence, predicting sequences of visual tokens that are subsequently transformed into high-quality video frames through diffusion rendering. This method is akin to traditional movie production processes, where complex stories are factorized down into manageable scene capturing. Further, we employ a multimodal script that enriches scene descriptions with detailed character information and visual style, enhancing continuity and character identity across scenes. We present extensive experiments across various movie genres, demonstrating that our approach not only achieves superior visual and narrative quality but also effectively extends the duration of generated content significantly beyond current capabilities.
@@ 1 Introduction
Driven by the advent of generative modeling techniques (rombach2022high; podell2023sdxl), there has been significant progress in video generation in the research community (guo2023animatediff; ho2022video; blattmann2023stable). Tremendous research efforts have focused on adapting a pre-trained text-to-image diffusion model (guo2023animatediff; ho2022video) to a video generation model. Recently revolutionary leap has been made with the Sora (videoworldsimulators2024) model which substantially scales a spatial-temporal transformer model. This model demonstrates remarkable video generation quality, dramatically extending the duration of generated videos from seconds to an unprecedented full minute. This technology profoundly reshapes the boundaries of generative AI, fostering anticipation towards hours-long movie production.
While predominant video generation methods adopt diffusion models, this paradigm is more suited for visual rendering and is less adept at modeling complex abstract logic and reasoning compared to autoregressive models, as evidenced in natural language processing. Moreover, diffusion models lack the flexibility to support arbitrary length. On the other hand, autoregressive models (wang2023neural; jiang2022vima; hu2023gaia1; pan2023kosmosg) have shown superior ability in handling complex reasoning and are better at predicting what might happen in the next—a key element of a world model. Moreover, autoregressive models offer greater flexibility in handling varied lengths of predictions and can also benefit from mature training and inference infrastructure. As such, there are attempts that use the autoregressive model for video generation (kondratyuk2023videopoet; hu2023gaia1). However, autoregressive modeling is not as compute efficient as diffusion model for visual rendering and demands substantially more computing even for image rendering.
@@
Recent progress (henschel2024streamingt2v; chen2023seine; qiu2023freenoise; videoworldsimulators2024) suggests that video generation quality in short duration (from seconds to a minute) can be consistently improved by scaling the compute due to the scaling law. In this work, we sidestep the short video generation problem and instead seek a generative framework for long video generation featuring complex narrative structures and intricate plot progressions, which is hardly achieved by merely scaling the current approaches. Our aim is to push the boundaries of video generation beyond the limitations of short-duration content, enabling the creation of videos with rich, engaging storylines.
@@
To this end, we represent the whole movie using sparse keyframes and employ a diffusion autoencoder to tokenize each keyframe into compact visual tokens. We train an autoregressive model to predict the sequence of visual tokens. This model is initialized from a pretrained large language mode, a 7B LLaMA model (touvron2023llama), to leverage its rich world knowledge for better generalizability. As such, the model can be regarded as a multimodal model, which, rather than outputting the text, predicts the visual tokens with movie script conditioning.
@@ 2 Related Work
Video generation has advanced mainly through diffusion and autoregression methods. The popularization of Stable Diffusion (rombach2022high) has sparked extensive studies, improving video quality with various techniques (guo2023animatediff; ho2022video; ho2022imagen; he2023animateastory). For long videos, strategies like (yin2023nuwaxl; qiu2023freenoise; wang2023genlvideo; chen2023seine; henschel2024streamingt2v) can only generate simple results. Despite their high computational load and challenges in managing complex sequences, diffusion models are widely used. Autoregression, on the other hand, is less common but excels in specific areas like autonomous driving by employing unified token spaces (hu2023gaia1), although it struggles with continuity in complex scenes. Our approach combines the strengths of both methods in a hierarchical and scalable manner, enhancing control and flexibility irrespective of video length.
@@
Recent advancements in NLP lead to the emergence of visual language models (VLMs) through enabling large language model (touvron2023llama) to comprehend visual modalities (liu2023llava; bai2023qwen; bai2023qwen2; chen2024internvl). In addition to language, VLMs have demonstrated their power in generating images (pan2023kosmosg; sun2023generative; Emu2). Moreover, ongoing research (zhang2023videollama; lin2023videollava; li2023llamavid; zhao2022lavila) has showcased the robust capabilities of Multimodal Language Models (MLLMs) in comprehending video content. Some works (cai2024spatialbot; cheng2024spatialrgpt; ren2024pixellm; rasheed2024glamm; lai2024lisa) enable VLMs to assist in spatial understanding and other visual tasks. VLMs (jiang2022vima; kim2024openvla; etukuru2024robotutilitymodelsgeneral; brohan2023rt; driess2023palm) also demonstrate the powerful capabilities in robotics. These advancements have solidified the effectiveness of VLMs in addressing complex problems. Building on this foundation, our approach leverages VLMs to facilitate the generation of extremely long video content.
@@
The crux of visual story generation lies in ensuring consistency across generated images. Similar to video generation, diffusion and autoregression are also two primary manners. Existing diffusion-based methods predominantly employ conditional generation to maintain the coherence  (su2023makeastoryboard; wang2023autostory; tewel2024trainingfree; hertz2023StyleAligned). Some (li2023photomaker; wu2024infiniteid) have shown promising results in retaining high-quality facial features and identity. Other approaches are based on the autoregressive ideas  (feng2023improved; liu2024intelligent; Rahman_2023_CVPR; pan2022synthesizing). However, these approaches generally yield mediocre results and are challenging to scale up. Story and video generation share intrinsic links, with overlapping solutions. Our method uniquely integrates these, offering a unified approach.
@@ 3.2 Keyframe Tokenization via Diffusion Autoencoder
[…]
@@
where $\mathbf{x}_{0}$ is the input image and $\mathbf{z}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\bm{\epsilon}$ is its corresponding noisy latent at timestep $t$. $\bm{\epsilon}\in\mathcal{N}(0,I)$ is the Gaussian noise. $\alpha_{t}$ and $\sigma_{t}$ define the noise schedule. In our experiment, we empirically find that merely two tokens can sufficiently characterize major semantics of keyframes, corroborating the finding of previous work (yang2022paint).
@@ 3.3 ID-preserving Diffusion Rendering
In our approach, while the primary diffusion decoder $\mathcal{D}$ adeptly reconstructs target images, it occasionally falls short in capturing fine-grained details, notably in facial features, due to the attenuation of details in compressed tokens. To address this, we enhance the cross-attention module within $\mathcal{D}$. This enhancement involves the integration of both descriptive text embedding, $\mathbf{d}$, and face embedding, $\mathbf{f}$, derived from the multimodal script. Specifically, the face embedding $\mathbf{f}$ and description embedding $\mathbf{d}$ is projected to the same dimension as the compressed token with two additional MLPs, which are then concatenated together and serve as the input to the keys and values of cross-attention modules in $\mathcal{D}$. The training loss is:
$\mathcal{L}_{DAE}=\mathbb{E}_{\mathbf{x}_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I})}\|\epsilon-\mathcal{D}(\mathbf{z}_{t},t,\mathcal{E}(\mathbf{x}_{0}),\mathbf{f},\mathbf{d})\|^{2}_{2},$ (2)
To further advocate the model’s ability to focus on critical details, we introduce a random masking strategy that obscures a subset of the input tokens as zero tokens only during training. This technique encourages the decoder to more effectively utilize the available facial and textual cues to reconstruct the images with higher fidelity, particularly in preserving identity-specific characteristics. This ID-preserving rendering also compensates for the loss of identity during the autoregressive modeling, leading to substantially improved identity perception quality as illustrated in Figure 3.
@@ 3.4 Autoregressive Keyframe Token Generation
We initiate our approach by leveraging a pretrained large language model, specifically LLaMA2-7B (touvron2023llama), to construct our autoregressive model, $\mathcal{G}$. Unlike traditional LLMs, $\mathcal{G}$ is designed to predict compressed vision tokens from the multimodal script $\mathbf{M}$ of the corresponding keyframe and historical information $\mathbf{H}$, which is formulated as $\mathbf{e}_{t}=\mathcal{G}(\mathbf{H}_{<t},\mathbf{M}_{t})$.
Traditional LLMs often employ cross-entropy loss for training, which is suitable for discrete outputs. However, our model deals with continuous real-valued image tokens, which makes cross-entropy inapplicable. Inspired by  tschannen2024givt, we adopt a $k$-mixture Gaussian Mixture Model (GMM) to effectively model the distribution of these real-valued tokens. This involves parameterizing the GMM with $kd$ means, $kd$ variances, and $k$ mixing coefficients.These parameters are predicted by attaching an additional trainable linear layer at the end of the autoregressive model, enabling the construction of the GMM and the sampling of continuous tokens from the GMM. The model is trained by minimizing the negative log-likelihood:
@@
We develop a well-structured multi-modal script format to serve as input for autoregressive models, as depicted in Figure 16. Our script integrates multiple dimensions: characters, scene elements, and plots. Accurately representing character appearance using text alone is challenging; therefore, we combine textual descriptions with facial embeddings (zheng2022general) to provide a more detailed representation of each character. To facilitate the processing by the autoregressive model, we structure the script format to distinctly delineate these elements.
@@
This method significantly extends the usable contextual length during training by condensing an entire sentence into a single token. We utilize LongCLIP (zhang2024longclip) as our text encoder for descriptions, supporting inputs up to 248 tokens, which enhances our capacity to handle detailed narrative content. Consequently, the multi-modal script at timestep $t$ and preceding historical information are represented as:
@@
To promote personalized movie content generation, we propose a few-shot learning method that utilizes in-context learning. During training, we select 10 random frames from an episode, encode them into visual tokens, and prepend these tokens stochastically to the input. This strategy not only promotes in-context learning, allowing the model to tailor content based on the reference frames, but also acts as a data augmentation technique, effectively mitigating overfitting.
@@ 3.5 Keyframe-based Video Generation
After obtaining the keyframes, we can generate video clips of the movie based on these keyframes. A straightforward approach is to utilize an existing image-to-video model, such as Stable Video Diffusion (SVD) (blattmann2023stable), to generate these clips. Specifically, SVD transforms the input image into latent features for conditioning and introduces interaction with the CLIP features of the input image via cross-attention. Although SVD is capable of generating high-quality short videos, e.g., 25 frames, it struggles to generate longer movie clips.
@@ 4.1 Implementation Details
We use pretrained models to reduce the training computational overhead, which is very common in the era of large models. We use the U-Net of pretrained SDXL (podell2023sdxl) as our diffusion decoder. The encoder consists of a pretrained CLIP image encoder (ViT-L) (radford2021learning) and a transformer-based compressor. The autoregressive model is initialized with a pretrained LLaMA-7B (touvron2023llama) checkpoint. All MLPs used to align different modal inputs follow the architecture in LLaVA (liu2023llava). Please refer to Appendix A for more details about the network architecture and training.
@@ 4.2 Experimental Setup
The majority of our training data is sourced from Condensed Movies (DBLP:conf/accv/BainNBZ20), with additional data collected using the same methodology as this dataset, resulting in 5M keyframes with corresponding script annotations. To systematically evaluate the effectiveness of our method, we construct a test dataset consisting of 100 long movies that are NOT included in the training set, with 1M keyframes after pre-processing. Each of these videos is processed through our proposed data pre-processing pipeline (see Appendix C), which accurately extracts and annotates keyframes.
@@
We evaluate the effectiveness of various methods using the following metrics. The CLIP score (radford2021learning) assesses semantic alignment between the plot and generated outputs. Quality is evaluated using the Aesthetic Score (AS) (schuhmann2022laion5bopenlargescaledataset), Fréchet Image Distance (FID) (heusel2018ganstrainedtimescaleupdate), and Inception Score (IS) (salimans2016improvedtechniquestraininggans). For assessing consistency, we propose new metrics for short-term (ST) and long-term (LT) consistency. ST is measured by calculating the CLIP cosine similarity between adjacent frames containing the target character. LT consistency is evaluated by checking the presence of the target character in selected frames across the test set. Finally, we calculate the ratio of frames containing the target character to the total number of selected frames as a measure of the accuracy in maintaining long-term character consistency. Please refer to the Appendix B for more details about the evaluation metrics.